

********************************************************************
UARS WEEKLY STATUS REPORT - 10 JUN 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM resumed operations on June 7. Since then, SUSIM has
successfully gathered its daily mid (1.1 nm) resolution, MgII, and
continuous monitoring scans for time periods with UARS telemetry
or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005039 has been processed and is waiting for archival
at the DAAC. HEPSA data for 2005040 to 2005049 is currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005039 has been processed and is waiting for DAAC archival.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 28 May 2005 (Orbit 75046, GMT Day 148) through
03 June 2005 (Orbit 75151, GMT Day 154).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-NONE THIS REPORT PERIOD.
A Reverse-to-Forward Yaw Around maneuver (number 140) was successfully
executed in Orbit 75072 beginning 149/ 18:22:34Z and completing at
149/ 18:49:44Z on 29 May 2005. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 265 degrees, the effective night length of the spacecraft
after the maneuver was approximately 53.25 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 25.76 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 24.35 and 12.23 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 10
July 2005 (DOY 191).
The beta angle ranged between the angles of 05.9 degrees to 18.9
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 19 June 2005 (DOY 170).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 6.13 to 6.09 degrees C. Additional
battery performance data for each battery on 27 May 2005 and 03 June
2005 are provided in Appendix D.
The clock error ranged from 10.8 to -9.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the status of the
batteries. The batteries maintained their VT-6 charge level
throughout the low beta/yaw period. The first scheduled turn on after
the yaw will be HALOE on 04 June 2005.
The FOT reviewed the latest tests performed on the solar array drive
in order to devise some new test scenarios. Four more ideas surfaced
for future testing, however this testing is now secondary to the UARS
OBC clock rollover effort. The clock rollover effort is important
because the onboard computer's clock will rollover on 16 September
2005. If this were allowed to happen the spacecraft clock would think
it was 1997 and all time related processing would be erroneous.
Therefore, the FOT, with the help of the FSW group, will prepare an
uplink procedure to reset the epoch year onboard to prevent this from
happening. This procedure is almost complete and will undergo
simulator testing this week. The uplink to the spacecraft is
scheduled for 02 August 2005.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 July 2005 (DOY 191). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 03 June 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
75061 149/0057 05-175 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
75063 149/0420 05-176 OBC/CLK CLK ADJ FR 20.6 TO -20.6 MSEC/DAY
75072 149/1756 05-177 MACS YAW - REV TO FWD
75088 150/1959 05-178 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
75150 154/2244 05-179 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 03 June 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 03 June 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
75061 149/0123 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
75063 149/0428 CLK ADJ FR 20.6 TO -20.6 MSEC/DAY
75088 150/1959 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
75150 154/2244 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 03 June 2005
27 May 2005 - GMT Day 147 Beta = 08.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 16.8/ -05.6 -4.31/-5.26 +00.0/+00.0 05.8 0.00 0.0
2 11.2/-123.2 +7.08/+5.79 +33.2/-18.8 25.3 1.03 23.5
3 72.8/-700.0 +1.97/+1.34 +19.6/-12.0 25.3 1.05 12.6
03 June 2005 - GMT Day 154 Beta = 18.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 11.2/ -22.4 -3.68/-4.94 +00.0/+00.0 05.8 0.00 0.0
2 151.2/ -89.6 +8.73/+7.74 +35.2/-18.4 26.1 1.04 23.4
3 -22.4/-700.0 +2.60/+1.65 +17.6/-09.6 26.1 1.08 10.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 03 June 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT 75114 152/1233 30832 Generic Late Acquisition
05-031 FOT 75122 153/0111 39664 Late Acquisition
05-030 FOT 75132 153/1734 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-028 FOT See List May 05 30832 Generic Late Acquisition
05-029 FOT 74949 141/1423 39569 NO ODMS
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT See List June 05 30832 Generic Late Acquisition
05-031 FOT 75122 153/0111 39664 Late Acquisition